OpenStudy (arbyj):

16. x + y = -3 how do i find the slope?

OpenStudy (anonymous):

rise over sun or sun over rise which one is it?

OpenStudy (anonymous):

or y=mx+b

OpenStudy (anonymous):

b is the y intercept and m is the slope which is u have to find it

OpenStudy (anonymous):

so the answer is -1

OpenStudy (unanimoose):

@katragaddasaichandra What? Did you even convert it yet?

OpenStudy (anonymous):

I did

OpenStudy (unanimoose):

Ahhhh my bad, yes it is -1.

OpenStudy (anonymous):

it is -1

OpenStudy (anonymous):

yes

OpenStudy (unanimoose):

Because after subtracting x on both sides it is -x, which is the same as -1x which is our mx.

OpenStudy (anonymous):

u r right

OpenStudy (unanimoose):

My bad, I apologize.

OpenStudy (arbyj):

can one of you guys show me a demonstration

OpenStudy (unanimoose):

Let's say we had: 3x + y = 6 We have to isolate y, so we move 3x to the other side while also removing it from the left. We do this by subtracting 3x from both sides so: 3x - 3x + y = -3x + 6 (-3x + 6 is the same as 6 - 3x, but we are doing it this way to go with y = mx + b). Now we have: y = -3x + 6 So our slope is -3 and our y intercept is 6.

OpenStudy (arbyj):

thank you